$SPAD/src/input card.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool card.output
)set message test on
)set message auto off
)clear all
--S 1 of 20
c0 := 0 :: CardinalNumber
--R
--R
--R   (1)  0
--R                                                         Type: CardinalNumber
--E 1
--S 2 of 20
c1 := 1 :: CardinalNumber
--R
--R
--R   (2)  1
--R                                                         Type: CardinalNumber
--E 2
--S 3 of 20
c2 := 2 :: CardinalNumber
--R
--R
--R   (3)  2
--R                                                         Type: CardinalNumber
--E 3
--S 4 of 20
c3 := 3 :: CardinalNumber
--R
--R
--R   (4)  3
--R                                                         Type: CardinalNumber
--E 4
--S 5 of 20
A0 := Aleph 0

--R
--R
--R   (5)  Aleph(0)
--R                                                         Type: CardinalNumber
--E 5
--S 6 of 20
A1 := Aleph 1
--R
--R
--R   (6)  Aleph(1)
--R                                                         Type: CardinalNumber
--E 6
--S 7 of 20
finite? c2
--R
--R
--R   (7)  true
--R                                                                Type: Boolean
--E 7
--S 8 of 20
finite? A0
--R
--R
--R   (8)  false
--R                                                                Type: Boolean
--E 8
--S 9 of 20
countable? c2
--R
--R
--R   (9)  true
--R                                                                Type: Boolean
--E 9
--S 10 of 20
countable? A0
--R
--R
--R   (10)  true
--R                                                                Type: Boolean
--E 10
--S 11 of 20
countable? A1
--R
--R

--R   (11)  false
--R                                                                Type: Boolean
--E 11
--S 12 of 20
[c2 + c2, c2 + A1]
--R
--R
--R   (12)  [4,Aleph(1)]
--R                                                   Type: List(CardinalNumber)
--E 12
--S 13 of 20
[c0*c2, c1*c2, c2*c2, c0*A1, c1*A1, c2*A1, A0*A1]
--R
--R
--R   (13)  [0,2,4,0,Aleph(1),Aleph(1),Aleph(1)]
--R                                                   Type: List(CardinalNumber)
--E 13
--S 14 of 20
[c2**c0, c2**c1, c2**c2, A1**c0, A1**c1, A1**c2]
--R
--R
--R   (14)  [1,2,4,1,Aleph(1),Aleph(1)]
--R                                                   Type: List(CardinalNumber)
--E 14
--S 15 of 20
[c2-c1, c2-c2, c2-c3, A1-c2, A1-A0, A1-A1]
--R
--R
--R   (15)  [1,0,"failed",Aleph(1),Aleph(1),"failed"]
--R                                   Type: List(Union(CardinalNumber,"failed"))
--E 15
--S 16 of 20
generalizedContinuumHypothesisAssumed true
--R
--R
--R   (16)  true
--R                                                                Type: Boolean
--E 16
--S 17 of 20
[c0**A0, c1**A0, c2**A0, A0**A0, A0**A1, A1**A0, A1**A1]
--R
--R
--R   (17)  [0,1,Aleph(1),Aleph(1),Aleph(2),Aleph(1),Aleph(2)]
--R                                                   Type: List(CardinalNumber)

--E 17
--S 18 of 20
a := Aleph 0
--R
--R
--R   (18)  Aleph(0)
--R                                                         Type: CardinalNumber
--E 18
--S 19 of 20
c := 2**a
--R
--R
--R   (19)  Aleph(1)
--R                                                         Type: CardinalNumber
--E 19
--S 20 of 20
f := 2**c
--R
--R
--R   (20)  Aleph(2)
--R                                                         Type: CardinalNumber
--E 20
)spool
)lisp (bye)

———-

References^{[1] nothing}

7